Grade 11IIT JEE Entrance Exam

the displacement y of a body varies with time t as y=-3/2t2+36t=2.how long does body takes to come to rest?

Nakshi Shah
8 Years agoGrade 11

1 Answer

Pavan sai vardhineedi
8 Years ago
It makes a quadratic equation when solved. 
36t-3t2/2=2
=3t2+4/2=36t
3t2-72t+4=0
But the thing is it has no real roots!!!